OpenStudy (anonymous):

Alg 1 help! Will fan and give a medal! Simplify. -12x^4/x^4+8x^5? a)-12/1+8x; where x does not equal -1/8 b) -12/1+8x; where x does not equal -1/8,0 c)-12/9x; where x does not equal 0 d) -12/9 **

OpenStudy (anonymous):

Would you want to help me? @haleyelizabeth2017 ?

OpenStudy (haleyelizabeth2017):

I can try :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (haleyelizabeth2017):

Shoot! I haven't done this for 2 years.......

OpenStudy (anonymous):

I have other questions that I want to know if I got the answer right with..

OpenStudy (anonymous):

Would that be better maybe? They all pretty much look like this though.

OpenStudy (haleyelizabeth2017):

Oh....don't know if I can help with this....sorry!

OpenStudy (haleyelizabeth2017):

Plus I have to go to work.....good luck though!

OpenStudy (anonymous):

Thanks anyway..

OpenStudy (haleyelizabeth2017):

Sorry!

OpenStudy (anonymous):

@zoomiemoo , can you please help me?

OpenStudy (anonymous):

@whydude81 , would you help too? pls <3

OpenStudy (anonymous):

is it -12x^4/(x^4+8x^5)

OpenStudy (anonymous):

To solve this we can start by breaking it up

OpenStudy (anonymous):

-12x^4/x^4 = -12

OpenStudy (anonymous):

how do we break it up?

OpenStudy (anonymous):

-12x^4/8x^5 = -3/2x

OpenStudy (anonymous):

-3/(2x)

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

no. now we have to put the parts back together (-12 + -3/(2x)) or (-12/ + -12/(8x) this equals -12/1+8x

OpenStudy (anonymous):

How did you do that? I have a lot of problems like this one.

OpenStudy (anonymous):

additionally anything that will make the the bottom of the fraction 0 will not be a solution so x cannot equal 1/8 however 0 can be a solution because the bottom still equals 1 when x = 0

OpenStudy (anonymous):

I moment I have another student how needs assistance

OpenStudy (anonymous):

I am back

OpenStudy (anonymous):

let me right it out give me a moment

OpenStudy (anonymous):

Okay, thank you. So is it -12/1+8x where x does not equal -1/8 or is it -12/1+8x where x does not equal -1/8,0 ?

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

x can equal 0 because the denominator will equal 1 which will give you a solution answer is A

OpenStudy (anonymous):

Would you still like me to explain the problem again

OpenStudy (anonymous):

Yes, please :)

OpenStudy (anonymous):

I just want to know how to simplify rational expressions..

OpenStudy (anonymous):

This is another problem, Simplify. x^2-3x-18/x+3 ? How would I begin to solve this?

OpenStudy (anonymous):

hello